# Know Your Limits: A Survey of Abstention in Large Language Models

### v2

@@ 3.3.3 Output-processing Approaches
Directly asking LLMs to evaluate if their responses are certain or safe (usually in a different conversation) and to abstain if they are not has proven effective in improving LLM abstention behavior (Phute et al., 2024; Kadavath et al., 2022; Varshney et al., 2023; Ren et al., 2023). Chen et al. (2023b) use Soft Prompt Tuning to learn self-evaluation parameters for various tasks. Similarly, Wang et al. (2024e) prompt the LLM to append a [harmful] or [harmless] tag to each of its responses; however, this approach may encourage over-abstention (Varshney et al., 2023).
@@
Multi-LLM systems are effective in producing better overall responses, including improved abstention behavior. In 2-LLM systems, a test LLM is employed to examine the output of the first LLM and helps with abstaining. For example, Mielke et al. (2022) use a second LLM to adjust the linguistic confidence of a response based on the model’s confidence score. In Wang et al. (2024b), the test LLM is used to guess the most likely harmful query from the output and abstains if a harmful query is detected. Pisano et al. (2023) critique and correct a model’s original compliant response using a secondary LLM. Multi-LLM systems beyond two LLMs leverage different LLMs as experts to compete or cooperate to reach a final abstention decision (Feng et al., 2024b; Chen et al., 2023a). Zeng et al. (2024) employ a group of LLMs in a system with an intention analyzer, original prompt analyzer, and judge.
[…]
@@ 4.1 Evaluation Benchmarks
Here are datasets designed to measure whether LLM outputs are “safe,” i.e., align with widely held ethical values; these datasets may consist of prompts that are either inherently unsafe or likely to elicit unsafe responses from LLMs. Some datasets focus on specific aspects of safety. A main concern is toxicity, when models generate harmful, offensive, or inappropriate content. For instance, RealToxicityPrompts (Gehman et al., 2020) gathers prompts to study toxic language generation, while ToxiGen (Hartvigsen et al., 2022) and LatentHatred (ElSherief et al., 2021) address implicit toxic speech, and ToxicChat (Lin et al., 2023) collects data from real-world user-AI interactions. Beyond toxicity, Beavertails (Ji et al., 2023a) balances safety and helpfulness in QA, CValues (Xu et al., 2023a) assesses safety and responsibility, and Xstest (Röttger et al., 2024a) examines exaggerated safety behaviors. LatentJailbreak (Qiu et al., 2023) and Do-Anything-Now (Shen et al., 2023) collect augmented unsafe prompts for malicious purposes.
Comprehensive safety benchmarks attempt to encompass a range of concerns. Röttger et al. (2024b) conduct the first systematic review of open datasets for evaluating LLM safety. Do-Not-Answer (Wang et al., 2024c) includes instructions covering information hazards, malicious uses, and discrimination. XSafety Wang et al. (2023a) provides a multilingual benchmark covering 14 safety issues across 10 languages. SALAD-Bench (Li et al., 2024a) is a large-scale dataset with a three-tier taxonomy, evaluating LLM safety and attack-defense methods. SORRY-Bench (Xie et al., 2024) proposes a more fine-grained taxonomy and diverse instructions. Most relevant to abstention, WildGuard (Han et al., 2024) evaluates model refusal performance as a necessary component for safety.
[…]
@@ 4.2 Evaluation metrics
  • Abstain Estimated Calibration Error (Abstain ECE) (Feng et al., 2024b) modifies traditional ECE (Guo et al., 2017) by including abstention. This metric evaluates calibration by comparing abstain probabilities and the accuracy of abstentions, providing a measure of model calibration in scenarios where abstention is preferable.
@@
  • Area Under Risk-Coverage Curve (AURCC) (Si et al., 2023) computes, for any given threshold, an associated coverage and error rate (risk), which is averaged over all thresholds. Lower AURCC indicates better selective QA performance.
  • Area Under Accuracy-Coverage Curve (AUACC) (Cole et al., 2023; Yoshikawa and Okazaki, 2023; Xin et al., 2021) computes, for any given threshold, an associated coverage and accuracy, which is averaged over all thresholds. Higher AUACC indicates better performance.
@@
[…]
@@
Many works implement LLM-as-a-judge for abstention evaluation (Mazeika et al., 2024; Souly et al., 2024; Chao et al., 2024). Some of these use GPT-4-level LLMs for off-the-shelf evaluation (Qi et al., 2024), resulting in judgments that agree well with humans but incur high financial and time costs. Others explore supplementary techniques to boost the accuracy of the LLM judge such as (i) Chain-of-thought prompting: asking the LLM to “think step-by-step” before deciding whether to not answer (Qi et al., 2024; Xie et al., 2024); (ii) In-context-learning: using refusal annotations from a training set as in-context examples (Xie et al., 2024); or (iii) Fine-tuning LLMs for abstention evaluation (Huang et al., 2024a; Li et al., 2024a).
[…]
@@
Human evaluation for abstention focuses on understanding user perceptions of different abstention expressions and the relation to the usefulness of a model’s response. Wester et al. (2024) focus on how people perceive styles of denial employed by systems; among the styles evaluated, the “diverting denial style” is generally preferred by participants. Kim et al. (2024b) investigate how expressing uncertainty affects user trust and task performance, finding that first-person uncertainty phrases like “I’m not sure, but…” reduce users’ confidence in the system’s reliability and their acceptance of its responses.
[…]
@@ 5 Other Considerations for Abstention
Over-abstention occurs when models abstain unnecessarily, e.g., Varshney et al. (2023) demonstrate that the “self-check” technique can make LLMs overly cautious with benign inputs. Others similarly observe that instruction tuning with excessive focus on abstention can lead models to inappropriately refuse to respond (Cheng et al., 2024; Bianchi et al., 2024; Wallace et al., 2024; Brahman et al., 2024). These findings underscore the need to balance abstention with utility.
@@
Abstention is highly sensitive to prompt wording. Zhou et al. (2023a) show that high-certainty expressions in prompts result in accuracy drop compared to low-certainty expressions. Moreover, safety-driven abstention mechanisms are notably susceptible to manipulation. Studies show that social engineering techniques such as persuasive language and strategic prompt engineering can bypass established safety protocols (Xu et al., 2023b; Chao et al., 2023). Even ostensibly benign strategies like finetuning with safe datasets or modifying decoding algorithms can inadvertently undermine the safety alignment of LLMs (Qi et al., 2024; Huang et al., 2024a). Advanced manipulation tactics include persona-based attacks (Shah et al., 2023), cipher-based communications (Yu et al., 2023), and the translation of inputs into low-resource languages (Yong et al., 2023; Feng et al., 2024a). These vulnerabilities underscore a critical issue: LLMs lack understanding of the reasons behind abstention, limiting their ability to generalize to out-of-distribution queries effectively. Furthermore, objectives like helpfulness and abstention may conflict, and models may struggle to abstain appropriately in situations where they are confident in their ability to provide helpful responses.
@@
Abstention should not be viewed as the termination of a conversation, but rather as a step towards subsequent information acquisition. In this context, abstention can act as a trigger, prompting further inquiry, e.g., asking the user for more information or retrieving additional relevant data (Feng et al., 2024b; Li et al., 2024b). After abstaining, systems should seek out more information when appropriate, transforming abstention from a static endpoint into a dynamic, constructive component of dialogue progression.
@@
Users have different preferences for model abstention (Wester et al., 2024) based on individual differences and task-specific needs, and no one-size-fits-all solution exists (Kirk et al., 2023b). Personalized abstention mechanisms in LLMs will allow the model to dynamically adjust its abstention behavior based on a user’s profile, tolerance for conservative responses, interaction history, and specific requirements.
@@ 6 Conclusion
There are many under-explored and promising research directions in abstention, some of which are described in this survey. While prior work investigates abstention in specific QA settings, we encourage study of abstention as a meta-capability across tasks, as well as more generalizable evaluation and customization of abstention capabilities to user needs. Other important directions include: enhancing privacy and copyright protections through abstention-aware designs to prevent the extraction of personal private information and copyrighted text fragments; generalizing the concept of abstention beyond LLMs to vision, vision-language, and generative machine learning applications; and improving multilingual abstention, as significant performance discrepancies exist between high-resource and low-resource languages, necessitating further research to ensure consistent performance across different languages.

### v3

@@ 3.3.3 Output-processing Approaches
Chen et al. (2023b) use Soft Prompt Tuning to learn self-evaluation parameters for various tasks. However, directly asking LLMs to evaluate if their responses are certain or safe (usually in a different conversation), and to abstain if they are not, has proven effective in improving LLM abstention (Phute et al., 2024; Kadavath et al., 2022; Varshney et al., 2023; Ren et al., 2023; Feng et al., 2024b). Kim et al. (2024a) allow the LLM to iteratively provide feedback on its own responses and refine its answers; this method achieves improvements in safety even in non-safety-aligned LLMs. Wang et al. (2024d) enable LLMs to self-evaluate responses and append a [harmful] or [harmless] tag to each response; however, this approach may encourage over-abstention.
@@
Multi-LLM systems are effective in producing better overall responses, including improved abstention behavior. In 2-LLM systems, a test LLM is employed to examine the output of the first LLM and helps with abstaining. In Wang et al. (2024b), the test LLM is used to guess the most likely harmful query from the output and abstains if a harmful query is detected. Pisano et al. (2023) critique and correct a model’s original compliant response using a secondary LLM.
Multi-LLM systems beyond two LLMs leverage different LLMs as experts to compete or cooperate to reach a final abstention decision (Feng et al., 2024b; Chen et al., 2023a). As an example, Zeng et al. (2024) employ a group of LLMs in a system with an intention analyzer, original prompt analyzer, and judge.
Self Evaluation & LLM Collaboration Summary LLMs can struggle with self-evaluation, but combining multiple LLMs has been effective in enhancing abstention capabilities.
@@ 4.1 Evaluation Benchmarks
Here are datasets designed to measure whether LLM outputs are “safe,” i.e., align with widely held ethical values; these datasets may consist of prompts that are either inherently unsafe or likely to elicit unsafe responses from LLMs. Some datasets focus on specific aspects of safety. A main concern is toxicity, when models generate harmful, offensive, or inappropriate content. For instance, RealToxicityPrompts (Gehman et al., 2020) gathers prompts to study toxic language generation, while ToxiGen (Hartvigsen et al., 2022) and LatentHatred (ElSherief et al., 2021) address implicit toxic speech, and ToxicChat (Lin et al., 2023) collects data from real-world user-AI interactions. Beyond toxicity, Beavertails (Ji et al., 2023a) balances safety and helpfulness in QA, CValues (Xu et al., 2023a) assesses safety and responsibility, and Xstest (Röttger et al., 2024a) examines exaggerated safety behaviors. LatentJailbreak (Qiu et al., 2023) introduces a benchmark that assesses both the safety and robustness of LLMs. Do-Anything-Now (Shen et al., 2023) collects a set of unsafe prompts for malicious purposes.
Comprehensive safety benchmarks attempt to encompass a range of concerns. Röttger et al. (2024b) conduct the first systematic review of open datasets for evaluating LLM safety. Do-Not-Answer (Wang et al., 2024c) includes instructions covering information hazards, malicious uses, discrimination, exclusion and toxicity, misinformation harms, and human-computer interaction harms. XSafety Wang et al. (2023a) provides a multilingual benchmark covering 14 safety issues across 10 languages. SALAD-Bench (Li et al., 2024a) is a large-scale dataset with a three-tier taxonomy, evaluating LLM safety and attack-defense methods. SORRY-Bench (Xie et al., 2024) proposes a more fine-grained taxonomy and diverse instructions. Most relevant to abstention, WildGuard (Han et al., 2024) evaluates model refusal performance as a necessary component for safety.
Evaluation Benchmarks Summary Existing benchmarks primarily focus on a single perspective. Researchers could benefit from more comprehensive benchmarks encompassing examples across the query, model, and human values perspectives, that are capable of system-wide assessment.
@@ 4.2 Evaluation metrics
  • Abstain Estimated Calibration Error (Abstain ECE) (Feng et al., 2024b) modifies traditional ECE (Guo et al., 2017a) by including abstention. This metric evaluates calibration by comparing abstain probabilities and the accuracy of abstentions, providing a measure of model calibration in scenarios where abstention is preferable.
@@
  • Area Under Risk-Coverage Curve (AURCC) (Si et al., 2023; Yoshikawa and Okazaki, 2023) computes, for any given threshold, an associated coverage and error rate (risk), which is averaged over all thresholds. Lower AURCC indicates better selective QA performance.
  • Area Under Accuracy-Coverage Curve (AUACC) (Cole et al., 2023; Xin et al., 2021) computes, for any given threshold, an associated coverage and accuracy, which is averaged over all thresholds. Higher AUACC indicates better performance.
@@
Statistical Automated Evaluation Summary Overall performance: No single metric captures all aspects of performance. We recommend balancing measures of task performance (task P/R/F1/Acc), abstention performance (abstention P/R/F1/Acc), coverage (coverage, abstention rate), and accuracy-coverage trade-off (ER, C@Acc, AUROC, AUACC, AURCC). Error rates: To assess abstention-related error rates, URUP measures the false answering rate and ARSP the false abstention rate. [Uncaptioned image] Evaluating partial abstention Existing statistical automated evaluation methods focus on identifying and evaluating full abstention, but partial abstention can be incorporated as a weighted sum of whether the response includes abstention and/or answers. By weighting abstention and answer accuracy, we can better qualify a model’s actual performance.
@@
Many works implement LLM-as-a-judge for abstention evaluation (Mazeika et al., 2024; Souly et al., 2024; Chao et al., 2024). Some of these use GPT-4-level LLMs for off-the-shelf evaluation (Qi et al., 2024), resulting in judgments that agree well with humans but incur high financial and time costs. Others explore supplementary techniques to boost the accuracy of the LLM judge such as (i) Chain-of-thought prompting: asking the LLM to “think step-by-step” before deciding whether to not answer (Qi et al., 2024; Xie et al., 2024); (ii) In-context-learning: using refusal annotations from a training set as in-context examples (Xie et al., 2024); or (iii) Finetuning LLMs for abstention evaluation (Huang et al., 2024a; Li et al., 2024a). Röttger et al. (2024a) extended full abstention evaluation by prompting GPT-4 with a taxonomy to classify responses as full compliance, full refusal, or partial refusal in a zero-shot setting.
Model-based Evaluation Summary Model-based evaluations focus on the human values perspective, particularly safety. Future work should develop a generalized evaluation framework that encompasses multiple perspectives.
@@
Human evaluation for abstention focuses on understanding user perceptions of different abstention expressions and the relation to the usefulness of a model’s response. Instead of binary decisions (full compliance and full refusal), Röttger et al. (2024a) introduce partial refusal when manually annotating model’s response. Wester et al. (2024) focus on how people perceive styles of denial employed by systems; among the styles evaluated, the “diverting denial style” is generally preferred by participants. Kim et al. (2024b) investigate how expressing uncertainty affects user trust and task performance, finding that first-person uncertainty phrases like “I’m not sure, but…” reduce users’ confidence in the system’s reliability and their acceptance of its responses.
Human-centric Evaluation Summary Human evaluation methods focus on full abstention and strong abstention expression categories, overlooking the significance of partial abstention and other forms of abstention expressions. Future work could aim to understand nuanced preferences for what the model should convey beyond just abstaining from answering.
@@ 5 Other Considerations for Abstention
Over-abstention occurs when models abstain unnecessarily. For example, Varshney et al. (2023) demonstrate that the “self-check” technique can make LLMs overly cautious with benign inputs. Others similarly observe that instruction tuning with excessive focus on abstention can lead models to inappropriately refuse to respond (Cheng et al., 2024; Bianchi et al., 2024; Wallace et al., 2024; Brahman et al., 2024). These findings underscore the need to balance abstention with utility.
@@
Abstention is highly sensitive to prompt wording. Safety-driven abstention mechanisms are notably susceptible to manipulation. Studies show that social engineering techniques such as persuasive language and strategic prompt engineering can bypass established safety protocols (Xu et al., 2023b; Chao et al., 2023). Even ostensibly benign approaches like finetuning with safe datasets or modifying decoding algorithms can inadvertently undermine the safety alignment of LLMs (Qi et al., 2024; Huang et al., 2024a). Advanced manipulation tactics include persona-based attacks (Shah et al., 2023), cipher-based communications (Yuan et al., 2024a), and the translation of inputs into low-resource languages (Yong et al., 2023; Feng et al., 2024a). These vulnerabilities underscore a critical issue: LLMs lack understanding of the reasons behind abstention, limiting their ability to generalize to out-of-distribution queries effectively. Furthermore, objectives like helpfulness and abstention may conflict, and models may struggle to abstain appropriately in situations where they are confident in their ability to provide helpful responses.
@@
Abstention should not be viewed as the termination of a conversation, but rather as a step towards subsequent information acquisition. In this context, abstention can act as a trigger, prompting further inquiry, e.g., asking the user for more information or retrieving additional relevant data (Feng et al., 2024b; Li et al., 2024b). After abstaining, systems should seek out more information when appropriate, transforming abstention from a static endpoint into a dynamic, constructive component of dialogue progression. For example, Zhao et al. (2024a) study the alternate task of reformulating unanswerable questions to questions that can be answered by a given document.
@@
Users have different preferences for model abstention (Wester et al., 2024) based on individual differences Zhang et al. (2024b) and task-specific needs, and no one-size-fits-all solution exists (Kirk et al., 2023b). Personalized abstention mechanisms in LLMs will allow the model to dynamically adjust its abstention behavior based on a user’s profile, tolerance for conservative responses, interaction history, specific query needs, and any other requirements.
6 Future Directions
There are many under-explored and promising research directions in abstention, some of which are described in this survey. While prior work has explicitly investigated abstention in specific tasks or implicitly contributed to improved abstention behaviors, we encourage study of abstention as a meta-capability across tasks, as well as more generalizable evaluation and customization of abstention capabilities to user needs. Beyond what has been discussed previously, other important directions include: (i) enhancing privacy and copyright protections through abstention-aware designs to prevent the extraction of personal private information and copyrighted text fragments; (ii) generalizing the concept of abstention beyond LLMs to vision, vision-language, and generative machine learning applications; and (ii) improving multilingual abstention, as significant performance discrepancies exist between high-resource and low-resource languages, necessitating further research to ensure consistent performance across different languages.
@@ 7 Conclusion
Acknowledgements
This research was supported in part by the National Science Foundation under CAREER Grant No. IIS2142739, and by the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the authors and do not reflect the official policy or position of the Department of Defense or the U.S. Government. We also gratefully acknowledge support from the UW iSchool Strategic Research Fund and the University of Washington Population Health Initiative. We thank the authors of the cited papers for reviewing our descriptions of their work.